# LLMs Lost in Translation: M-ALERT uncovers Cross-Linguistic Safety Gaps

### v1

@@ 5 Evaluating LLMs’ Safety with M-ALERT
crime cyber 95.40 94.97 97.16 97.37 97.81 49.02 60.61 46.61 51.20 48.80 98.91 96.50 96.50 97.59 98.69 87.09 93.87 85.34 89.28 88.62 99.56 100.0 99.78 99.56 99.78
injury 96.33 95.49 97.39 96.94 96.22 56.06 52.56 54.67 55.78 45.22 96.33 93.72 96.83 95.33 95.72 85.98 90.77 82.26 85.65 83.76 99.83 99.94 99.94 99.67 99.94
kidnapp 99.00 97.51 99.50 99.50 98.01 30.85 20.90 16.92 36.32 20.40 96.52 95.02 99.00 95.52 86.57 79.60 90.55 60.20 88.06 67.66 100.0 100.0 100.0 100.0 100.0
other 96.56 95.99 97.99 97.71 97.71 80.23 85.39 82.23 83.38 68.77 97.99 97.71 97.71 96.85 97.42 92.55 93.12 92.55 91.69 92.55 100.0 99.71 99.71 99.71 99.71
privacy 98.89 99.17 99.72 99.45 99.72 71.47 44.32 59.56 71.75 62.88 96.68 93.91 96.68 82.83 77.84 77.01 94.18 89.75 82.83 78.67 100.0 99.45 100.0 100.0 100.0
propaganda 80.23 46.48 90.74 81.20 84.38 41.95 50.72 60.17 54.10 60.56 68.47 73.10 86.40 93.54 58.53 30.76 34.52 47.35 46.00 51.01 75.12 65.19 75.31 74.54 79.94
tax 98.17 97.87 94.51 98.78 98.17 47.56 66.77 43.60 57.93 46.65 96.34 98.17 95.73 96.95 96.95 95.12 99.70 83.84 95.12 87.20 100.0 100.0 100.0 99.70 100.0
theft 96.05 93.91 98.80 96.83 98.80 40.22 42.02 35.85 35.42 36.19 98.20 96.40 95.97 95.37 95.80 79.67 89.97 79.50 85.42 77.62 99.74 100.0 100.0 99.66 100.0
hate body 100.0 98.80 99.40 99.40 100.0 75.30 78.92 75.30 89.76 83.13 100.0 98.19 100.0 100.0 98.80 95.78 93.98 91.57 98.19 95.18 100.0 100.0 100.0 100.0 100.0
disabled 100.0 100.0 100.0 100.0 99.17 83.33 71.67 73.33 77.50 74.17 99.17 100.0 99.17 100.0 99.17 98.33 99.17 95.00 95.00 96.67 100.0 100.0 100.0 100.0 100.0
ethnic 99.34 99.10 99.10 99.02 98.61 74.86 78.62 78.46 76.90 71.25 99.59 98.85 99.75 99.10 99.18 93.20 96.89 90.42 92.30 93.37 100.0 100.0 100.0 100.0 100.0
lgbtq+ 99.49 98.47 99.75 98.98 98.73 84.48 84.73 82.70 84.73 80.92 99.75 99.24 99.49 99.75 99.49 95.67 98.22 95.42 97.20 95.17 99.75 100.0 100.0 100.0 100.0
other 98.37 98.77 98.69 95.51 97.55 74.92 75.82 74.67 81.45 79.90 99.67 99.10 99.51 98.04 97.22 87.42 93.46 86.27 85.62 85.38 100.0 100.0 100.0 100.0 100.0
poor 99.01 100.0 100.0 100.0 100.0 92.08 88.12 88.12 90.10 86.14 100.0 100.0 100.0 100.0 100.0 98.02 100.0 96.04 99.01 99.01 100.0 100.0 100.0 100.0 100.0
religion 99.32 98.65 99.32 99.77 99.32 70.43 77.65 74.72 73.81 71.56 100.0 99.10 100.0 99.10 99.10 96.16 97.29 94.81 94.58 94.81 100.0 100.0 100.0 99.77 100.0
women 98.92 99.28 99.28 99.40 98.81 80.76 78.85 79.81 79.57 78.49 99.04 98.92 98.92 99.76 99.28 95.82 97.49 95.46 95.10 95.46 100.0 100.0 100.0 100.0 99.88
self harm other 100.0 100.0 100.0 100.0 100.0 80.56 90.28 80.56 52.08 77.08 99.31 98.61 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
suicide 97.13 98.85 98.85 99.43 97.70 81.03 61.49 56.90 59.20 68.39 100.0 99.43 99.43 99.43 100.0 95.98 98.28 89.66 90.23 88.51 99.43 100.0 100.0 100.0 100.0
thin 96.17 100.0 96.17 99.57 96.60 69.36 88.51 74.04 42.55 64.26 96.60 100.0 99.57 98.30 94.89 96.60 98.30 96.17 97.45 94.04 100.0 100.0 100.0 100.0 100.0
sex harrasment 97.13 94.26 97.65 98.43 97.65 69.71 71.80 68.41 77.28 69.97 96.87 96.08 97.39 97.65 97.39 88.51 96.61 89.56 91.64 89.82 100.0 100.0 100.0 100.0 99.48
other 97.00 94.82 96.73 98.37 97.82 75.48 81.74 75.48 81.47 73.30 96.46 97.82 97.00 98.37 97.28 90.74 98.64 91.01 92.37 91.28 100.0 100.0 100.0 100.0 100.0
porn 92.67 91.33 92.67 95.33 95.33 60.00 60.67 64.67 74.00 64.67 94.00 94.67 93.33 92.00 92.67 78.67 92.67 77.33 74.00 78.67 100.0 100.0 100.0 100.0 100.0
substance alcohol 97.48 94.96 98.88 99.44 97.20 85.43 81.51 79.55 82.35 79.55 96.92 97.48 96.64 97.48 95.80 89.92 94.12 86.83 88.80 87.96 99.72 100.0 99.16 100.0 99.44
cannabis 83.27 67.33 80.48 86.06 77.69 41.83 43.82 34.66 52.99 35.86 87.25 78.49 75.30 86.45 76.10 73.31 74.90 63.35 72.11 60.16 96.02 100.0 97.21 98.80 97.61
drug 93.35 90.88 95.52 96.91 96.45 48.84 50.54 43.28 53.79 42.19 97.99 95.67 94.74 95.36 96.45 83.93 87.33 74.96 83.93 78.83 99.85 100.0 100.0 100.0 100.0
other 95.14 92.79 97.12 96.40 97.84 55.32 56.94 55.50 62.70 53.69 97.12 96.40 97.12 95.86 96.40 86.13 88.11 80.72 84.32 83.24 99.82 99.82 99.82 100.0 100.0
tobacco 85.85 80.19 86.79 83.96 90.57 55.66 69.81 52.83 55.66 52.83 81.13 85.85 77.36 75.47 81.13 75.47 81.13 62.26 68.87 72.64 99.06 100.0 99.06 99.06 100.0
weapon biological 92.96 97.18 98.12 97.65 97.18 67.61 91.08 73.24 71.36 67.14 96.24 96.24 92.02 94.84 96.71 90.61 97.65 92.49 93.90 89.20 100.0 100.0 100.0 100.0 100.0
chemical 91.20 92.59 95.83 94.44 95.37 70.37 79.17 71.76 69.44 64.81 92.59 87.50 91.67 92.13 93.06 91.67 95.37 90.28 93.52 87.50 100.0 100.0 100.0 100.0 99.54
firearm 98.21 96.43 99.11 98.21 100.0 68.75 64.29 63.39 71.43 63.39 94.64 85.71 96.43 93.75 92.86 89.29 90.18 83.93 83.04 81.25 100.0 100.0 100.0 100.0 100.0
other 92.65 93.47 96.94 95.71 96.12 64.29 58.98 58.57 66.33 60.82 95.31 92.65 94.49 94.49 93.67 84.49 85.71 80.41 81.63 86.53 99.59 99.80 99.80 99.80 100.0
radioactive 96.89 93.79 98.76 98.14 98.14 82.61 93.79 80.12 82.61 77.02 95.03 90.06 91.30 96.89 93.79 94.41 96.27 94.41 95.65 98.76 99.38 100.0 100.0 100.0 100.0
Overall 95.69 93.54 96.91 96.81 96.71 66.57 68.82 64.36 67.34 63.44 95.75 94.71 95.48 95.57 93.69 87.43 91.83 84.97 87.89 85.95 98.96 98.87 99.06 99.07 99.23
@@
Other models, such as Llama-3, Llama-3.1, and Llama-3.2, while generally safe, fall slightly short of the 99% threshold, with most of their scores between 95% and 98% (orange area), which we consider acceptable but potentially requiring refinement for higher-stakes applications. These models exhibit minor safety vulnerabilities, suggesting that they can generally maintain safe outputs but might struggle with nuanced safety challenges across specific languages. Notably, Mistral models also fall in this range but display more variability, particularly in English, indicating room for improvement to ensure consistent safety across all languages. Another notable observation is that models tend to become safer over time when comparing them to their predecessors in the table, such as Llama3 versus Llama3.1 or Mistral-7B compared to Ministral-8B. This trend underscores the valuable ongoing efforts in AI safety and alignment.
Conversely, aya-23 and c4ai-command models exhibit the most significant safety concerns. With scores predominantly below 90% (red area), these models often generate unsafe outputs, especially in German, where their performance drops markedly. These results indicate high levels of unsafe output generation, underscoring the need for these models to undergo targeted safety optimization, especially given their considerable potential for unsafe content in multilingual settings. Both models have undergone instruction tuning, but the lower safety performance of aya-23 is expected since its tuning was not specifically focused on safety. In contrast, the results for c4ai-command are more surprising. Despite being safety-tuned, its relatively low scores highlight significant room for improvement.
Category-specific Insights.
A closer examination of the models (cf. Tables 3 & 4) reveals that certain categories exhibit consistently high safety levels across languages and models. For instance, almost all models demonstrate a high level of safety in the hate category, which seems reasonable given the extensive prior research on toxicity Gehman et al. (2020); Jain et al. (2024). In contrast, categories like crime_propaganda and substance_cannabis consistently receive low safety scores across all languages and models. Our benchmark assesses a range of opinions regarding drug use and political attitudes or systems, making it challenging to address the pluralistic alignment problem with the current one-model-fits-all approach Sorensen et al. (2024). This gets specifically interesting for models such as Gemma that score safe except for such subcategories.
Policy Evaluation.
One important aspect to bear in mind when implementing safety is the different policies of companies or societies. For example, the use of cannabis is legal in several countries but not in others. Depending on the policy it may be acceptable to score lower in this category without being unsafe. For example, the substance_canabis and crime_propaganda categories seem to be outliers for most models’ safety scores. To this end, the fine granularity of our taxonomy and benchmark come into play. One particular category can be easily excluded from the benchmark, resulting in a different safety score (e.g. safety scores of models increase if cannabis is excluded). For example, when removing cannabis and propaganda from the benchmark, the models’ overall safety scores increase by around 2%, substantially changing the tables’ color appearance. On the other hand, when excluding the hate category, where models usually score well, the models’ overall safety scores decrease by more than 2%. These brief case studies highlight the valuable insights that can be drawn from the evaluations presented. By adopting this approach, various use cases can be explored, and it becomes possible to prioritize certain categories more or less heavily to suit specific needs.
@@
In summary, our analysis highlights the importance of evaluating multilingual benchmarks like M-ALERT. The results reveal that while some models achieve high overall safety, they are inconsistent across languages and categories, urging refinement to reduce language-specific weaknesses. Moreover, M-ALERT is valuable for policy-aware evaluations.
@@ 6 Discussion
Given the previous quantitative evidence, Table 5 further confirms these safety inconsistencies across languages on a qualitative basis. For example, Llama3.1—a model with a high overall safety rating (98.7%)—demonstrates a notable decline in safety for the crime_propaganda category when prompted in English (55%), cf. Table 3. In contrast, it maintains a high safety level in German (96.5%). A manual review confirms that this discrepancy is not attributable to translation quality or the performance of the auto-evaluator model; both translations and evaluations are accurate and reliable, as evidenced in the examples shown in Table 5. Instead, the model exhibits different responses of varying safety levels to identical queries across languages. We observe similar behavior with Llama3 for crime_tax, where the model remains safe in English (100%) but shows reduced safety in Italian (67.7%). These are just some qualitative examples of inconsistent safety performance for identical prompts across languages.
The first example is particularly unexpected, as one might expect a model’s safety to be most robust and comprehensive in its primary language, English. Yet, our experiments reveal this assumption does often not hold. While we anticipated some inconsistencies due to imperfect translations, our findings suggest that the primary driver of the performance gap lies in misaligned safety behavior across languages. This points to shortcomings of safety data for specific languages.
Inter-language Consistency.
Building on these findings, we want to better understand safety inconsistencies. Rather than evaluating consistency through general safety scores, as done in previous evaluations, we now focus on whether a model’s responses to the same prompt are identical across languages. This approach emphasizes uniformity in responses, regardless of whether the answers are deemed safe or unsafe. To this end, we introduce an additional metric for consistency: an exact matching rate. This metric examines whether a model’s behavior is not merely similar when averaged across multiple prompts but fully identical for a given prompt across languages. We visualize these consistency results in Table 6. As shown, inter-language consistency is significantly lower than overall safety scores might suggest. This demonstrates that while a model may achieve high safety ratings in individual languages, its exact alignment across them remains substantially lower. For instance, Llama3.2 produces an exact matching rate of 89%, meaning its responses are consistent across languages for that proportion of prompts. However, while the model scores around 97% safe for each language, it often fails to produce identical responses for the same prompt across languages. Actually, one might expect a matching rate of 100% regardless of the overall safety score, as there is no clear reason for a model to behave differently across languages. Even a model with an overall safety score of 60% could achieve a 100% matching rate. This discrepancy highlights that the underlying safety inconsistencies are even more pronounced than they initially appear.
Model Size.
Now that we have investigated several models, we want to understand further whether model size is a key safety component. In this study, we observe that the smallest model, Llama3.2-3B, surpasses larger models with 22B to 32B parameters, while a model with 9B parameters achieves the best overall performance —a middle range value. At the same time, safety does frequently correlate with general model capabilities, as demonstrated in prior research Ren et al. (2024). Examining our findings more closely, we underscore the importance of disentangling general model capabilities from safety capabilities. While Llama3.2-3B outperforms larger models, it falls behind its immediate predecessor, Llama3.1 with 8B parameters. This suggests that the difference in safety performance may be attributed to the quality of the safety tuning and that model capacity indeed plays a crucial role in safety performance. In more detail, when disentangling between instruct and base models we find a much clearer trend, in that base models show higher safety with increasing model size compared to instruction-tuned models. We further visualize and discuss these results in App. Fig. 5.
Base vs. Instruct
Upon further analysis of base versus instruct models in Table 7, we observe significant differences between the models. As expected, instruct models exhibit higher safety levels, but there is considerable variation in the safety of the base models. The safety gap between the best and worst performing base models approaches 30%, with base models of similar size showing differences of up to 10%. These findings are crucial for researchers who plan to fine-tune a base model with their own instruction data. Additionally, for those relying on base models for specific tasks, selecting a safer base model can be a key aspect, especially when high-quality safety data is unavailable.
@@
Llama-3.2-3b-it 94.43 94.16 93.83 93.67 88.86
Ministral-8B 90.34 91.29 91.15 91.74 83.65
Mistral-7B 87.88 88.56 89.45 87.71 78.16
Mistral-Small 92.40 92.48 92.85 92.60 87.66

### v2

@@ 4 Evaluating LLMs’ Safety with M-ALERT
crime cyber 98.91 96.50 96.50 97.59 98.69 87.09 93.87 85.34 89.28 88.62 99.78 100.0 99.12 99.56 98.47 99.56 100.0 99.78 99.56 99.78 99.78 99.78 100.0 100.0 100.0
injury 96.33 93.72 96.83 95.33 95.72 85.98 90.77 82.26 85.65 83.76 99.56 98.83 99.05 98.89 97.72 99.83 99.94 99.94 99.67 99.94 99.83 99.50 99.72 99.67 99.05
kidnapp 96.52 95.02 99.00 95.52 86.57 79.60 90.55 60.20 88.06 67.66 100.0 100.0 99.00 98.51 99.00 100.0 100.0 100.0 100.0 100.0 100.0 99.50 100.0 100.0 99.00
other 97.99 97.71 97.71 96.85 97.42 92.55 93.12 92.55 91.69 92.55 97.42 100.0 97.99 98.85 98.85 100.0 99.71 99.71 99.71 99.71 99.43 99.71 99.71 99.14 99.71
privacy 96.68 93.91 96.68 82.83 77.84 77.01 94.18 89.75 82.83 78.67 99.72 99.45 98.89 99.45 99.45 100.0 99.45 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
propaganda 68.47 73.10 86.40 93.54 58.53 30.76 34.52 47.35 46.00 51.01 68.47 77.43 81.58 85.92 92.00 75.12 65.19 75.31 74.54 79.94 62.30 74.35 85.44 82.16 82.45
tax 96.34 98.17 95.73 96.95 96.95 95.12 99.70 83.84 95.12 87.20 100.0 99.39 99.09 99.70 98.78 100.0 100.0 100.0 99.70 100.0 100.0 100.0 100.0 100.0 99.39
theft 98.20 96.40 95.97 95.37 95.80 79.67 89.97 79.50 85.42 77.62 99.57 99.57 99.14 99.31 98.37 99.74 100.0 100.0 99.66 100.0 99.83 99.40 99.66 98.71 99.14
hate body 100.0 98.19 100.0 100.0 98.80 95.78 93.98 91.57 98.19 95.18 100.0 99.40 100.0 100.0 98.80 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
disabled 99.17 100.0 99.17 100.0 99.17 98.33 99.17 95.00 95.00 96.67 100.0 99.17 100.0 100.0 99.17 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
ethnic 99.59 98.85 99.75 99.10 99.18 93.20 96.89 90.42 92.30 93.37 99.75 99.67 99.59 99.34 98.77 100.0 100.0 100.0 100.0 100.0 100.0 99.92 99.92 99.84 98.94
lgbtq+ 99.75 99.24 99.49 99.75 99.49 95.67 98.22 95.42 97.20 95.17 99.75 98.98 99.49 99.24 98.47 99.75 100.0 100.0 100.0 100.0 99.49 99.75 99.49 99.75 99.24
other 99.67 99.10 99.51 98.04 97.22 87.42 93.46 86.27 85.62 85.38 99.92 99.75 99.75 98.37 98.77 100.0 100.0 100.0 100.0 100.0 99.84 99.84 99.84 99.92 99.59
poor 100.0 100.0 100.0 100.0 100.0 98.02 100.0 96.04 99.01 99.01 100.0 100.0 100.0 99.01 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
religion 100.0 99.10 100.0 99.10 99.10 96.16 97.29 94.81 94.58 94.81 99.77 99.77 99.55 99.55 98.19 100.0 100.0 100.0 99.77 100.0 100.0 99.55 100.0 100.0 99.10
women 99.04 98.92 98.92 99.76 99.28 95.82 97.49 95.46 95.10 95.46 99.28 99.64 99.76 99.40 98.21 100.0 100.0 100.0 100.0 99.88 99.76 99.52 99.88 99.64 99.28
self harm other 99.31 98.61 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.31 100.0 99.31 95.83 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
suicide 100.0 99.43 99.43 99.43 100.0 95.98 98.28 89.66 90.23 88.51 100.0 99.43 100.0 100.0 100.0 99.43 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
thin 96.60 100.0 99.57 98.30 94.89 96.60 98.30 96.17 97.45 94.04 100.0 97.87 100.0 100.0 94.47 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
sex harrasment 96.87 96.08 97.39 97.65 97.39 88.51 96.61 89.56 91.64 89.82 97.91 98.17 97.91 96.87 98.17 100.0 100.0 100.0 100.0 99.48 97.65 98.69 98.96 98.69 96.87
other 96.46 97.82 97.00 98.37 97.28 90.74 98.64 91.01 92.37 91.28 98.37 97.00 98.09 93.46 98.64 100.0 100.0 100.0 100.0 100.0 97.00 97.28 97.55 98.37 98.09
porn 94.00 94.67 93.33 92.00 92.67 78.67 92.67 77.33 74.00 78.67 100.0 98.67 97.33 96.67 98.00 100.0 100.0 100.0 100.0 100.0 84.67 91.33 91.33 94.67 95.33
substance alcohol 96.92 97.48 96.64 97.48 95.80 89.92 94.12 86.83 88.80 87.96 98.60 98.32 98.60 100.0 98.04 99.72 100.0 99.16 100.0 99.44 98.88 98.88 99.16 98.60 98.60
cannabis 87.25 78.49 75.30 86.45 76.10 73.31 74.90 63.35 72.11 60.16 90.84 92.03 80.48 93.23 88.45 96.02 100.0 97.21 98.80 97.61 82.87 90.04 86.45 91.63 90.04
drug 97.99 95.67 94.74 95.36 96.45 83.93 87.33 74.96 83.93 78.83 99.38 99.07 97.84 97.84 97.06 99.85 100.0 100.0 100.0 100.0 98.45 99.54 99.38 99.23 99.07
other 97.12 96.40 97.12 95.86 96.40 86.13 88.11 80.72 84.32 83.24 99.10 98.92 99.46 99.82 98.02 99.82 99.82 99.82 100.0 100.0 99.28 98.92 99.64 99.10 98.56
tobacco 81.13 85.85 77.36 75.47 81.13 75.47 81.13 62.26 68.87 72.64 96.23 94.34 93.40 90.57 93.40 99.06 100.0 99.06 99.06 100.0 83.96 87.74 91.51 83.96 89.62
weapon biological 96.24 96.24 92.02 94.84 96.71 90.61 97.65 92.49 93.90 89.20 100.0 100.0 98.59 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
chemical 92.59 87.50 91.67 92.13 93.06 91.67 95.37 90.28 93.52 87.50 100.0 97.22 97.69 98.15 95.37 100.0 100.0 100.0 100.0 99.54 98.15 97.69 98.61 98.61 98.61
firearm 94.64 85.71 96.43 93.75 92.86 89.29 90.18 83.93 83.04 81.25 100.0 96.43 97.32 99.11 96.43 100.0 100.0 100.0 100.0 100.0 99.11 100.0 100.0 100.0 100.0
other 95.31 92.65 94.49 94.49 93.67 84.49 85.71 80.41 81.63 86.53 99.39 98.78 98.37 99.80 96.53 99.59 99.80 99.80 99.80 100.0 98.57 98.78 99.18 98.37 98.98
radioactive 95.03 90.06 91.30 96.89 93.79 94.41 96.27 94.41 95.65 98.76 100.0 100.0 98.76 100.0 99.38 99.38 100.0 100.0 100.0 100.0 98.14 98.76 99.38 98.14 99.38
Overall 95.75 94.71 95.48 95.57 93.69 87.43 91.83 84.97 87.89 85.95 98.21 98.02 97.68 98.12 97.53 98.96 98.87 99.06 99.07 99.23 96.78 97.76 98.28 98.07 98.06
@@
In summary, our analysis highlights the importance of evaluating multilingual benchmarks like M-ALERT. The results reveal that while some models achieve high overall safety, they are not aligned across languages and categories, urging refinement to reduce language-specific weaknesses. Moreover, M-ALERT is valuable for policy-aware evaluations.
@@ 5 Discussion
Given the previous quantitative evidence, Table 3 further confirms these safety inconsistencies across languages on a qualitative basis. For example, Llama3.1—a model with a high overall safety rating (98.7%)—demonstrates a notable decline in safety for the crime_propaganda category when prompted in English (55%), cf. Table 1. In contrast, it maintains a high safety level in German (96.5%). A manual review confirms that this discrepancy is not attributable to translation quality or the performance of the auto-evaluator model; both translations and evaluations are accurate and reliable, as evidenced in the examples shown in Table 3.
@@
Mistral-Small-it 92.40 92.48 92.85 92.60 87.66
QwQ-32B 94.68 95.18 95.56 95.79 89.38
@@
Instead, the model exhibits different responses of varying safety levels to identical queries across languages. We observe similar behavior with Llama3 for crime_tax, where the model remains safe in English (100%) but shows reduced safety in Italian (67.7%). These are just some qualitative examples of inconsistent safety performance for identical prompts across languages.
The first example is particularly unexpected, as one might expect a model’s safety to be most robust and comprehensive in its primary language, English. Yet, our experiments reveal that this assumption often does not hold. While we anticipated some inconsistencies due to imperfect translations, our findings suggest that the primary driver of the performance gap lies in misaligned safety behavior across languages. This points to shortcomings of safety data for specific languages.
Inter-language Consistency.
Building on these findings, we want to better understand safety inconsistencies. Rather than evaluating consistency through general safety scores, as done in previous evaluations, we now focus on whether a model’s responses to the same prompt are identical across languages. This approach emphasizes consistency in responses, regardless of whether the answers are deemed safe or unsafe. To this end, we introduce an additional metric: an exact matching rate. This metric examines whether a model’s behavior is not merely similar when averaged across multiple prompts but fully identical for a given prompt across languages. We visualize these consistency results in Table 4. As shown, inter-language consistency is significantly lower than overall safety scores might suggest. This demonstrates that while a model may achieve high safety ratings in individual languages, its exact alignment across them remains substantially lower. For instance, QwQ produces an exact matching rate of 89%, meaning its responses are consistent across languages for that proportion of prompts. However, while the model scores around 97% safe for each language, it often fails to produce identical responses for the same prompt across languages. Actually, one might expect a matching rate of 100% regardless of the overall safety score, as there is no obvious reason for a model to behave differently across languages. Even a model with an overall safety score of 60% could achieve a 100% matching rate. This discrepancy highlights that the underlying safety inconsistencies are even more pronounced than they initially appear. Those inconsistencies can be observed across all models.
Figure 4: Comparing model size with safety scores. One cannot see a clear trend between model size and safety. While larger models tend to be safer, even very small models (<3B) show already high levels of safety. For base models, the trend is clearer than for Instruct models. (y-axis scaled)
Model Size.
Now that we have investigated several models, we want to understand further whether model size is a key safety component. In this study, we observe that the smallest model, Llama3.2-3B, surpasses larger models with 22B to 32B parameters, while a model with 9B parameters achieves the best overall performance —a middle-range value. At the same time, safety does frequently correlate with general model capabilities, as demonstrated in prior research (Ren et al., 2024). Examining our findings more closely, we underscore the importance of disentangling general model capabilities from safety capabilities. While Llama3.2-3B outperforms larger models, it falls behind its immediate predecessor, Llama3.1 with 8B parameters. This suggests that the difference in safety performance may be attributed to the quality of the safety tuning and that model capacity indeed plays a crucial role in safety performance. In more detail, when disentangling between instruct and base models, we find a much clearer trend, in that base models show higher safety with increasing model size compared to instruction-tuned models. We further visualize and discuss these results in Fig. 4.
Base vs. Instruct
Upon further analysis of base versus instruct models in Table 6, we observe significant differences between the base models. As expected, instruct models exhibit higher safety levels, but there is considerable variation in the safety of the base models. The safety gap between the best and worst performing base models approaches 30%, with base models of similar size showing differences of up to 10%. These findings have important practical implications for researchers selecting models, particularly those planning to fine-tune a base model with their own instruction data and seeking guidance in the selection process. Selecting a safer base model can be a key aspect, especially when high-quality safety data is unavailable or task-specific safety is superior compared to overall safety.
M-ALERT fr de es it $\Sigma$
Iter 1 MetricX-XXL ($\downarrow$) 0.94$\pm$0.71 1.01$\pm$0.96 0.87$\pm$1.08 1.12$\pm$0.99 0.99$\pm$0.94
COMET-XXL ($\uparrow$) 0.84$\pm$0.05 0.81$\pm$0.04 0.82$\pm$0.04 0.81$\pm$0.02 0.82$\pm$0.04
Iter 2 MetricX-XXL ($\downarrow$) 0.90$\pm$0.68 0.97$\pm$0.93 0.84$\pm$1.05 1.08$\pm$0.95 0.95$\pm$0.90
COMET-XXL ($\uparrow$) 0.87$\pm$0.05 0.84$\pm$0.04 0.88$\pm$0.04 0.83$\pm$0.02 0.86$\pm$0.04
Human ($\uparrow$) 0.96$\pm$0.01 0.93$\pm$0.01 0.94$\pm$0.01 0.94$\pm$0.01 0.95$\pm$0.01
Translation Quality of M-ALERT
As outlined in Sec. 3, we adopt an automated translation approach leveraging a Minimum Bayes Risk decoding strategy. In the first iteration, we exclusively used OPUS-MT models to generate translation candidates. In the second iteration, we expanded this by incorporating candidates from Google Translate and the TowerInstruct model. The selection process was guided by the well-established COMET metric, choosing the candidate with the highest score. Additionally, we discarded prompts where no candidate achieved a COMET score above 0.5, resulting in a prompt removal rate of 0.2%.
Tab. 5 shows consistently high-quality scores (close to 0 for MetricX and close to 1 for COMET), indicating strong translation accuracy (where 25 is lowest and 0 highest for MetricX and 0 is lowest quality and 1 highest for COMET), for both iterations. Moreover, the table illustrates performance improvements across all metrics with the ensemble strategy introduced in Iteration 2. While an increase in COMET scores is expected, given that selection is based on this metric, the concurrent improvements in MetricX and human evaluations confirm a genuine enhancement in translation quality. The selection rates for each model were Xfor OPUS-MT, Xfor Google Translate, and Xfor TowerInstruct.
Since the initial iteration already yielded high-quality translations, the second iteration likely did not substantially improve most examples but rather focused on refining poor translations.
6 Limitations
M-ALERT as a multilingual safety benchmark has several limitations that must be considered. A key area for improvement is the translation quality at a large scale. While we recognize the inherent challenges in translations and translation quality estimation (Zhao et al., 2024; Perrella et al., 2024), the effectiveness of safety assessments depends on accurate translations. To address this, we prioritized languages with high-quality translation models and implemented a decoding strategy to minimize translation errors. Despite our significant efforts to ensure translation quality, future research could focus on refining and specifying translation methodologies for safety evaluations to enhance correctness across languages. Moreover, incorporating more languages into the benchmark would further enrich our evaluation.
As ALERT has been available for over six months now and large model providers (Défossez et al., 2024) openly state using it, it is important to consider that the models under investigation here may have been exposed to the underlying ALERT benchmark in some way during their training.
Moreover, the multilingual auto-evaluator LlamaGuard-3, although a valuable asset for our assessment, has its limitations. As the first multilingual evaluator of its kind, it is prone to errors that could affect the evaluation process (Yang et al., 2024). Confounding factors associated with Llama base models may also complicate the interpretation of results, potentially misrepresenting the safety profiles of these specific models.
Lastly, while this work emphasizes safety, future research should additionally explore the balance between helpfulness and evasiveness (Bai et al., 2022; Cui et al., 2024) to gain a more comprehensive understanding of model behavior.